Do the following equations represent parallel lines? Why (or why not)? 2y=3x+4 and 2y=3x-4
3y=4x-1 and 4y=-3x-1

Answer by richwmiller(17219)

The first two are parallel 2y=3x+4 2y=3x-4 because they have the same slope
3/2
the second two are perpendicular because the product of their slopes =-1
m=4/3 and m=-3/4